# Unlocking Efficient Long-to-Short LLM Reasoning with Model Merging

### v1

@@ 3.1 Model Merging
Ties Merging
TIES Merging (Yadav et al., 2023) is an efficient method for integrating parameters from multiple FT models, addressing redundancy and conflicts. Its key steps include: (1) pruning parameters, retaining significant deviations from pre-trained weights; (2) resolving conflicts via majority voting or alignment; and (3) weighted aggregation of significant parameters to form the final model. This approach reduces noise and enhances generalization, particularly for integrating fine-tuned large language models (LLMs) across related tasks.
@@ 6 Further Analysis
From the perspective of datasets, as shown in Table 6, longer responses are generated for more challenging datasets. This trend is consistent across the System 1 quick-thinking model, the System 2 reasoning model, and the merged models. Specifically, when examining the MATH500 dataset with varying difficulty levels, the same conclusion can be drawn: response length positively correlates with the difficulty level of the questions.
@@
Based on these observations, we hypothesize how the merged models maintain reasoning accuracy while reducing response length. For merged models with a high proportion of reflective responses, such as LoRE-Merging and Sens-Merging, the reduction in response length is primarily achieved by eliminating redundant reasoning content and streamlining the reasoning process, while still performing reflections in most cases. This assumption is supported by the statistics on the MATH500 dataset, which indicate that the frequency of reflection keywords appearing in each response of Sens-Merging and LoRE-Merging is 2.5, compared to 11.1 per response of the DeepSeek-R1-7B model. In contrast, merged models like Tie-Merging and AIM-Ties inherently possess the ability to quickly respond to simpler questions and engage in more deliberate reasoning for complex ones, as evidenced by the distribution of reflective responses across various datasets.
@@
Unlike the training-based methods (Ma et al., 2025; Xia et al., 2025) which need to change the base model for training, model merging offers a more straightforward approach to achieving short-to-long adjustments in quick-thinking models by simply tuning the merging weights or selecting an appropriate base merging model, without the need for further training. For instance, in a simple attempt to achieve short-to-long reasoning, we assigned a small weight (e.g., 0.2) to the slow-thinking model while using the quick-thinking model as the base. This approach resulted in an approximate improvement of over 20 points in the overall score, accompanied by a 25%+ increase in response length. model merging provides a more effective and efficient solution for enabling System 1 models to acquire System 2 reasoning abilities compared to model distillation (Yu et al., 2024b).
@@ A.1 Experiments Configuration
We report our merging configurations on various merging methods as Table 7. All merged models are saved and evaluated in BF16, except for those produced by Sens-Merging, as the sensitivity factors are calculated on a CPU, which does not support BF16 computations. To preserve the accuracy of the calculated sensitivity scores, the merged models generated through Sens-Merging are saved in FP32. For evaluations, we use the standard Qwen evaluation toolkit. For System 1 models, we use the ‘cot’ prompt with few-shots demonstrations and a maximum length of 8192. For System 2 models, we use ‘qwen25-math-cot’ prompt in a zero-shot setting with a maximum length of 10240. We also conducted experiments on LLaMA3.1-8B models; however, we were unable to reproduce the reasoning accuracy reported in their work using the Qwen evaluation toolkit. As a result, we do not report the results for the LLaMA-series models.

### v2

@@ 3.1 Model Merging
TIES-Merging
TIES-Merging (Yadav et al., 2023) is an efficient method for integrating parameters from multiple FT models, addressing redundancy and conflicts. Its key steps include: (1) pruning parameters, retaining significant deviations from pre-trained weights; (2) resolving conflicts via majority voting or alignment; and (3) weighted aggregation of significant parameters to form the final model. This approach reduces noise and enhances generalization, particularly for integrating fine-tuned large language models (LLMs) across related tasks.
@@ 6 Further Analysis
From the perspective of datasets, as shown in Table 1, longer responses are generated for more challenging datasets. This trend is consistent across the System 1 quick-thinking model, the System 2 reasoning model, and the merged models. Specifically, when examining the MATH500 dataset with varying difficulty levels, the same conclusion can be drawn: response length positively correlates with the difficulty level of the questions, as shown in Figure 2(a).
@@
Based on these observations, we hypothesize how the merged models maintain reasoning accuracy while reducing response length. For merged models with a high proportion of reflective responses, such as LoRE-Merging and Sens-Merging, the reduction in response length is primarily achieved by eliminating redundant reasoning content and streamlining the reasoning process, while still performing reflections in most cases. This assumption is supported by the statistics on the MATH500 dataset, which indicate that the frequency of reflection keywords appearing in each response of Sens-Merging and LoRE-Merging is 2.5, compared to 11.1 per response of the DeepSeek-R1-7B model. In contrast, merged models like TIES-Merging and AIM-TIES inherently possess the ability to quickly respond to simpler questions and engage in more deliberate reasoning for complex ones, as evidenced by the distribution of reflective responses across various datasets.
@@
Unlike the training-based methods (Ma et al., 2025; Xia et al., 2025) which need to change the base model for training, model merging offers a more straightforward approach to achieving short-to-long adjustments on quick-thinking models by simply tuning the merging weights or selecting an appropriate base merging model, without the need for further training. For instance, in a simple attempt to achieve short-to-long reasoning, we assigned a small weight (e.g., 0.2) to the slow-thinking model while using the quick-thinking model as the base. This approach resulted in an approximate improvement of over 20 points in the overall score, accompanied by a 25%+ increase in response length. Model merging provides a more effective and efficient solution for enabling System 1 models to acquire System 2 reasoning abilities compared to model distillation (Yu et al., 2024b).
@@ A.1 Experiments Configuration
We report our merging configurations on various merging methods as Table 7. All merged models are saved and evaluated in BF16, except for those produced by Sens-Merging, as the sensitivity factors are calculated on the CPUs, which does not support BF16 computations. To preserve the accuracy of the calculated sensitivity scores, the merged models generated through Sens-Merging are saved in FP32. For evaluations, we use the standard Qwen evaluation toolkit. For System 1 models, we use the ‘cot’ prompt with few-shots demonstrations and a maximum length of 8192. For System 2 models, we use ‘qwen25-math-cot’ prompt in a zero-shot setting with a maximum length of 10240. We also conducted experiments on LLaMA3.1-8B models; however, we were unable to reproduce the reasoning accuracy reported in their work using the Qwen evaluation toolkit. As a result, we do not report the results for the LLaMA-series models.